# Aligning MLLM Benchmark With Human Preferences via Structural Equation Modeling

### v1

@@ Abstract
Evaluating multimodal large language models (MLLMs) remains a fundamental challenge due to a lack of structured, interpretable, and theoretically grounded benchmark designs. Existing benchmarks often adopt heuristic-based task groupings with unclear cognitive targets, thus resulting in overlapping abilities, redundant indicators, and limited diagnostic power. In this work, we propose a novel framework for aligning MLLM benchmark based on Structural Equation Modeling (SEM) to analyze and quantify the internal validity, dimensional separability, and contribution of benchmark components. Motivated by the observed limitations of current designs, we further introduce a novel capability hierarchy grounded in Piaget’s theory of cognitive development, dividing MLLM abilities into three hierarchical layers, i.e., Perception, Memory, and Reasoning. We reorganize existing MLLM benchmarks under the proposed framework and construct a new benchmark named Gold. Experimental results demonstrate that the proposed benchmark exhibits stronger interpretability, reduced indicator redundancy, and clearer cognitive consistency compared to existing approaches. Our benchmark is available at https://github.com/tianyu-zou/GoldBench.
1 Introduction
@@
The rapid advancements in the field of multimodal learning have been driven by the emergence of increasingly Multimodal Large Language Models (MLLMs). From OpenAI’s GPT-4o achiam2023gpt and Google’s Gemini team2023gemini to a variety of open-source models such as LLaVA liu2023visual ; touvron2023llama , mPLUG-Owl ye2023mplug , and BLIP-2 li2023blip , these models exhibit remarkable capabilities in integrating and processing multiple modalities, including text, images, and videos. As MLLMs continue to evolve, the need for comprehensive evaluation frameworks becomes increasingly critical to assess their reasoning abilities, multimodal understanding, and generalization performance fu2024mme ; li2024embodied . A rigorous evaluation is essential not only for benchmarking advancements in model development but also for identifying potential biases, addressing ethical concerns, and ensuring the responsible and reliable deployment of these models across various real-world applications feng2025benchmarking ; qin2025survey ; zhang2024q .
@@
Evaluations not only reveal the advantages and shortcomings of these models across various tasks and dimensions, but they also provide clear directions for future enhancements zhu2023large ; li2024survey ; huang2024survey . Currently, the evaluation of MLLMs primarily relies on benchmark datasets composed of large-scale question-answer pairs guo2025r ; li2025benchmark . According to statistics from Opencompass 2023opencompass , there are now over 300 evaluation benchmarks for MLLMs, with more than 100 of them integrated into various multimodal evaluation suites. However, the rapid proliferation of evaluation benchmarks has also introduced new challenges: the lack of unified standards makes it difficult to assess the quality, coverage, and reliability of these benchmarks li2025information ; zhang2025redundancy . As a result, it becomes increasingly hard to determine which benchmarks offer meaningful insights into model capabilities and which may lead to biased or misleading conclusions.
@@
A particularly notable trend in recent benchmarks is the widespread adoption of tiered capability frameworks, that is attempts to deconstruct model performance into structured dimensions. We have found that while most evaluation frameworks embrace this approach, there exist significant differences among them in terms of capability dimensions and task design. Typical evaluation benchmarks include: SEED-Bench li2024seed , which proposes a five-level capability framework and constructs 27 tasks; MV-Bench li2024mvbench , which divides 47 evaluation tasks based on six core capabilities; and MMT-Bench ying2024mmt , which expands capability dimensions to 32 with an accompanying 162 fine-grained tasks. This phenomenon of ”dimension expansion” reflects the ongoing academic exploration of model understanding. However, it also raises a fundamental methodological concern: How can we construct a theoretically grounded and empirically verifiable capability structure that avoids arbitrary task proliferation and ensures interpretability?
To address this, we elaborate a novel framework by using Structural Equation Modeling (SEM) ullman2012structural as a principled approach to define and validate the latent structure of model abilities based on observed task performance. Recognizing human evaluation as the gold standard for reliability, its extensive labor cost and limited scalability create an urgent need for an alternative that approximates human judgment while dramatically reducing manual effort. The proposed framework meets this requirement by modeling the causal relationships between observed task indicators and the unobservable constructs they form. It organizes characteristics such as perception and cognition into empirically testable measurement models. Unlike traditional statistical or heuristic analyses, it not only infers latent variables from observable data but also reveals the causal contributions of individual features to overall evaluation quality, and can be calibrated to align with human cognitive preferences. By adopting the proposed framework for the evaluation of MLLM benchmarks, we can systematically assess the structural validity and practical effectiveness of the benchmark.
@@
To further leverage our evaluation framework, we draw on Piaget’s theory of cognitive development piaget2000piaget , using its account of how humans acquire perceptual, mnemonic, and reasoning abilities to derive the human-centered Gold benchmark. Traditional benchmarks, driven largely by heuristic task groupings, often yield inconsistent or overlapping ability hierarchies that neither reflect human cognition nor guide task design in a principled way. A Piaget-inspired bottom-up stratification offers a clear remedy. By aligning MLLM capabilities with the Sensorimotor, Preoperational, and (Concrete & Formal) Operational stages of human development, we define three continuous ability layers in Gold: Perception, Memory, and Reasoning. This human-cognitive foundation enhances interpretability and consistency in benchmarks.
@@
Our main contributions are as follows:
  • SEM-based Evaluation Framework: To the best of our knowledge, we for the first time apply SEM in MLLM assessment, which models the causal relationships between observable task metrics and latent capabilities. This approach not only introduces a novel evaluation pipeline but also ensures that it provides a transparent, verifiable, and reproducible framework for MLLM benchmark evaluation.
  • Piaget-Inspired Hierarchical Capability Structure: By leveraging Piaget’s theory of cognitive development, we define a three-tier capability hierarchy, i.e., Perception, Memory, and Reasoning, which helps us systematically organize tasks, mitigate dimension inflation, and enhance interpretability.
  • Gold Benchmark: We propose the human-centered Gold benchmark, derived from our Piaget-inspired hierarchical framework, to systematically evaluate MLLM across perceptual, mnemonic, and reasoning tasks in a structured, cognitively grounded manner.
@@
2 Evaluating Benchmark Structures via SEM
2.1 Revisiting of Structural Equation Modeling
SEM can be constructed based on causal measurement models, organizing multiple observed variables into a coherent structural framework. It consists of two core components: the measurement model, which defines the relationships between observed variables and their underlying latent constructs, and the structural model, which captures the causal relationships among latent variables. As discussed in Appendix, we adopt a formative measurement model as shown in Figure 1. The measurement model specifies the relationships between observed and latent variables as follows: $X=\Lambda_{x}\xi+\delta$, $Y=\Lambda_{y}\eta+\epsilon$, where $X$ and $Y$ are observed exogenous and endogenous variables, $\xi$ and $\eta$ are latent variables, $\Lambda_{x}$ and $\Lambda_{y}$ are factor loading matrices, and $\delta$, $\epsilon$ are measurement errors. The structural model describes causal relations among latent variables: $\eta=B\eta+\Gamma\xi+\zeta$, with $B$ and $\Gamma$ representing path coefficients and $\zeta$ the residuals. Together, the measurement and structural models provide a comprehensive framework for assessing both direct and indirect effects.
2.2 Structural Analysis of Existing Benchmarks
@@
Interestingly, we further apply the proposed SEM-based evaluating framework to analyze existing MLLM benchmarks. Specifically, we investigate whether the task configurations within current benchmarks provide complementary information while jointly forming meaningful latent constructs. Our analysis reveals that most benchmarks struggle to balance low redundancy with high construct validity.
@@
To illustrate this issue in detail, we conduct a case study on two representative benchmarks: MMBench liu2024mmbench and BLINK fu2024blink . Leveraging the performance reports of over 190 MLLMs collected via the VLMEvalKit duan2024vlmevalkit toolkit, we construct formative structural equation models based on the hierarchical task taxonomy specified by each benchmark. In our modeling framework, individual tasks serve as observed indicators, which are grouped into intermediate latent variables representing ability dimensions such as perception, reasoning, and cross-modal understanding. These first-order latent abilities are then aggregated into a second-order latent variable that captures the overall capability of an MLLM. Crucially, to align this latent construct with human-perceived performance, we incorporate the Chatbot Arena chiang2024chatbot score as an external observed indicator of the top-level latent variable. This hierarchical SEM structure enables us to evaluate how well benchmark-defined ability dimensions, as derived from task-level scores, explain human-judged overall performance.
@@
In the case of MMBench, as shown in Figure 2(a), tasks exhibit high contribution to their corresponding ability dimensions, with outer loadings exceeding 0.85 for most tasks. However, the analysis of Variance Inflation Factor (VIF) values in Figure 2(b) shows a concerning level of multicollinearity, particularly for tasks like Image Scene (VIF = 13.96) and Attribute Recognition (VIF = 6.12), indicating a high degree of redundancy within the benchmark. This redundancy undermines the overall efficiency of the measurement model, as the tasks are provide insufficiently independent information.
On the other hand, BLINK as shown in Figure 3 demonstrates a different issue. While the VIF values for BLINK are generally low (with a maximum value of 3.09 for Art Style), indicating that task redundancy is less of a concern, the outer loadings of several tasks are lower compared to MMBench. Tasks such as IQ Test (outer loading = 0.354) and Counting (outer loading = 0.686) contribute less to their respective constructs, suggesting that the benchmark lacks sufficiently strong task performance indicators for each cognitive ability dimension.
@@
Despite structural differences, both MMBench and BLINK exhibit limited discriminant validity among their ability dimensions, as indicated by their respective HTMT matrices. Several HTMT values exceed or approach the 0.90 threshold, particularly between conceptually adjacent dimensions, revealing substantial overlap and suggesting that current ability taxonomies in both benchmarks require further conceptual clarification.
Overall, while MMBench and BLINK differ in their structural characteristics, MMBench showing strong but redundant task contributions, and BLINK presenting low redundancy but weaker indicators. Both suffer from poor discriminant validity in their ability dimensions. These findings expose fundamental limitations in current benchmark design and underscore the need for a more systematic approach to evaluating benchmark structure. These findings reveal the urgent need for a principled framework to evaluate the internal structure of MLLM benchmarks and to redefine ability dimensions based on sound cognitive foundations.
2.3 A SEM-based Evaluation Framework for Benchmark
@@
Dimensional Diversity.
Dimensional diversity aims to quantify the discriminant validity among latent ability dimensions, thereby characterizing their conceptual independence and potential redundancy. Although existing benchmarks effectively capture multiple facets of MLLMs’ capabilities, overlaps between different dimensions may still exist. To rigorously assess discriminant validity, we adopt the Heterotrait–Monotrait Ratio of correlations (HTMT). For any two latent dimensions $\xi_{i}$ and $\xi_{j}$, the HTMT index is defined as
@@
$D_{\mathrm{div}}=\frac{1}{2*\max_{i\neq j}\mathrm{HTMT}_{ij}}$ (2)
where $\max{i\neq j}\mathrm{HTMT}{ij}$ represents the maximum HTMT value across all distinct pairs of latent dimensions. The denominator is scaled by 2 to ensure that $D{\mathrm{div}}\in[0,1]$, given that HTMT values are typically bounded within the empirical range of [0, 1]. A higher $D_{\mathrm{div}}$ indicates stronger discriminant validity and thus greater conceptual distinctiveness among latent constructs.
Task Contribution.
Task contribution quantifies the relative importance of each task indicator in shaping its corresponding latent construct. In order to capture this notion in a formative measurement context, we adopt the concept of out-loading, which represents the standardized weight of each observed task on the latent dimension. To evaluate the contribution of tasks at the level of the entire benchmark, which consists of multiple ability dimensions $\xi_{=}\{\xi_{1},\dots,\xi_{m}\}$, we define the benchmark-level task contribution score as
$\begin{split}TC&=\frac{1}{\sum_{i=1}^{n}m_{i}}\sum_{i=1}^{m_{i}}\sum_{j=1}^{n}\left|\lambda_{i,j}\right|,\end{split}$ (3)
the metric $TC\in[0,1]$ provides a global estimate of task-level informativeness across the entire benchmark, with higher values indicating stronger and more concentrated contributions from individual tasks.
Indicator Validity.
Indicator validity is assessed both substantively and statistically, with the latter focusing on multicollinearity diagnostics among task indicators. For each indicator $x_{i,j}$ associated with latent dimension $\xi_{j}$, we compute the Variance Inflation Factor (VIF) by
@@
where $R^{2}_{x_{i,j}}$ is the coefficient of determination obtained by regressing $x_{i,j}$ on the remaining tasks. A higher VIF value indicates stronger collinearity, suggesting that the task may not provide unique information beyond what is already captured by others.
To consolidate individual VIF values into a single benchmark-level measure, we define the composite validity score $D_{\mathrm{valid}}$ as the inverse of the geometric mean of all task-level VIFs

### v2

@@ Abstract
Evaluating multimodal large language models (MLLMs) remains a fundamental challenge due to a lack of structured, interpretable, and theoretically grounded benchmark designs. Existing benchmarks often adopt heuristic-based task groupings with unclear cognitive targets, thus resulting in overlapping abilities, redundant indicators, and limited diagnostic power. To do as, we propose a novel framework for aligning MLLM benchmark based on structural equation modeling to analyze and quantify internal validity, dimensional separability, and contribution of benchmark components. Motivated by the observed limitations of current designs, we further introduce a novel capability hierarchy grounded in Piaget’s theory of cognitive development, dividing MLLM abilities into three hierarchical layers, i.e., Perception, Memory, and Reasoning. We reorganize existing MLLM benchmarks under the proposed framework and construct a new benchmark named Gold. Experimental results demonstrate that the proposed benchmark exhibits stronger interpretability, reduced indicator redundancy, and clearer cognitive consistency compared to existing approaches. Our benchmark is available at https://github.com/tianyu-zou/GoldBench.
I Introduction
@@
The rapid advancements in the field of multimodal learning have been driven by the emergence of increasingly powerful and versatile Multimodal Large Language Models (MLLMs) [1, 2, 3]. From OpenAI’s GPT-4o [4] and Google’s Gemini [5] to a variety of open-source models such as LLaVA [6, 7], mPLUG-Owl [8], and BLIP-2 [9], these models exhibit remarkable capabilities in integrating and processing multiple modalities, including text, images, and videos. As MLLMs continue to evolve [10, 11], the need for comprehensive evaluation frameworks becomes increasingly critical to assess their reasoning abilities, multimodal understanding, and generalization performance [12, 13]. A rigorous evaluation is essential not only for benchmarking advancements in model development but also for identifying potential biases, addressing ethical concerns, and ensuring the responsible and reliable deployment of these models across various real-world applications [14, 15, 16].
Figure 1: Brief illustrations of How to Evaluate Benchmark?
@@
Evaluations not only reveal the advantages and shortcomings of these models across various tasks and dimensions, but they also provide clear directions for future enhancements [17, 18, 19]. Currently, the evaluation of MLLMs primarily relies on benchmark datasets composed of large-scale question-answer pairs [20, 21]. According to statistics from Opencompass [22], there are now over 300 evaluation benchmarks for MLLMs, with more than 100 of them integrated into various multimodal evaluation suites. However, the rapid proliferation of evaluation benchmarks has also introduced new challenges: the lack of unified standards makes it difficult to assess the quality, coverage, and reliability of these benchmarks [23, 24]. As a result, it becomes increasingly hard to determine which benchmarks offer meaningful insights into model capabilities and which may lead to biased or misleading conclusions. As shown in Fig. 1, we develop a structural equation modeling [25] based framework to evaluate the quality and contribution of MLLM benchmarks, where SEM serves as a statistical approach to link observed task performances with latent capability dimensions, enabling us to evaluate the quality and contribution of MLLM benchmarks.
TABLE I: Comprehensive Benchmark Overview
Benchmark
Capability
Dimension
Task
Number
Publication
Date
MME [26] 4 14 Jun-23
LVLM-eHub [27] 6 16 Jun-23
SEED-Bench [28] 2 12 Aug-23
TouchStone [29] 5 27 Sep-23
LAMM [30] 4 9 Nov-23
ChEF [31] 6 9 Nov-23
SEED-Bench-2 [32] 6 27 Nov-23
SEED-Bench-2-Plus [32] 3 63 Apr-24
MMT-Bench [33] 32 162 Apr-24
BLINK [34] 3 14 Jul-24
TinyLVLM [35] 6 42 Aug-24
UniBench [36] 17 53 Aug-24
MMStar [37] 6 18 Sep-24
MMBench [38] 6 20 Sep-24
MM-Vet [39] 6 16 Dec-24
@@
Recent representative multi-modal benchmarks have shown a trend toward the widespread adoption of tiered capability frameworks, which aim to deconstruct model performance into structured dimensions. As illustrated in Table I, there is a clear pattern of increasing the number of tasks and capability dimensions across recent benchmarks. This trend is particularly evident within the same benchmark series. For instance, SEED-Bench [32] initially proposes a two-dimensional framework with 12 tasks, but its updated version, SEED-Bench-2 [32], expands to six dimensions and 27 tasks. Similarly, other benchmarks also exhibit this “dimension expansion”, such as MVBench [40], which divides 47 tasks based on six core capabilities, and MMTBench [33], which introduces 32 dimensions and 162 fine-grained tasks. While this ongoing expansion reflects a deeper exploration of model understanding, it also raises an important methodological issue: How can we construct a theoretically grounded and empirically verifiable capability structure that avoids arbitrary task proliferation and ensures interpretability?
To address this issue, we develop a novel SEM-based pruning framework to define and validate the latent structure of model abilities based on observed task performance. While human evaluation is widely regarded as the gold standard for assessing reliability, its high labor cost and limited scalability make it impractical for large-scale benchmark analysis [41]. This creates an urgent need for an alternative that can approximate human judgment while substantially reducing manual effort. Guided by human evaluation results [42, 43, 44] as the reliability reference, our framework constructs SEM models to capture the latent structure of model abilities from observed task performance. In this process, SEM not only infers unobservable constructs such as perception and cognition, but also quantifies the causal contributions of individual tasks to the overall evaluation [45]. This enables the pruning of redundant tasks while retaining those most aligned with human judgment.
@@
Building on this SEM-based pruning framework, we further ground the construction of the benchmark in Piaget’s theory of cognitive development [46], using its account of how humans acquire perceptual, mnemonic, and reasoning abilities to derive the human-centered Gold benchmark. Traditional benchmarks, driven largely by heuristic task groupings, often yield inconsistent or overlapping ability hierarchies that neither reflect human cognition nor guide task design in a principled way [47]. A Piaget‑inspired bottom‑up stratification offers a clear remedy. By aligning MLLM capabilities with the Sensorimotor, Preoperational, and (Concrete & Formal) Operational stages of human development, we define three continuous ability layers in Gold: Perception, Memory, and Reasoning. This human‑cognitive foundation enhances interpretability and consistency in benchmarks.
Our main contributions are as follows:
  • SEM-based Evaluation Framework: To the best of our knowledge, we are the first to apply SEM in MLLM assessment, modeling the causal relationships between observable task metrics and latent capabilities. This approach not only introduces a novel evaluation pipeline but also ensures a transparent, verifiable, and reproducible framework for MLLM benchmark evaluation.
  • Piaget-Inspired Hierarchical Capability Structure: By leveraging Piaget’s theory of cognitive development, we define a three-level ability hierarchy, Perception, Memory, and Reasoning, which systematically organizes tasks, mitigates dimension inflation, and enhances interpretability.
  • Gold Benchmark: We propose the human-centered Gold benchmark, derived from our Piaget-inspired hierarchical framework, to evaluate MLLMs across perceptual, mnemonic, and reasoning tasks in a structured and cognitively grounded manner.
TABLE II: Multimodal Benchmarks by Capability
Capability Benchmarks
Perception Q-Bench [16], PEC [48], GVT-Bench [49], VBench [50], OCRBench [51], CODE [52], MMUBench [53], MMVP [54], CVBench [55], EQBEN [56], P2GB [57], MDVP-Bench [58], MM-SAP [59], MagnifierBench [60], UNIAA [61], AesBench [62], II-Bench [63], ImplicitAVE [64], EmoBench [65], FABA-Bench [66], LaVy-Bench [67], MMMB [68], M3GIA [69], SeaEval [70], CVQA [71], Henna [72], MTVQA [73]
Memory CODIS [74], MMNeedle [75], MileBench [76], MM-NIAH [77], MuirBench [78], Mementos [79], MMIU [80], Mantis-Eval [81], IIT [82], VEGA [83], MMC4 [84], Obelics [85], CoMM [86], VL-ICLBench [87], MMMU [88], VideoNIAH [89], OSCaR [90], TempCompass [91], VITATECS [92], MovieChat-1K [93], EgoSchema [94], TimeIT [95], ADLMCQ [96], MLVU [97], Event-Bench [98], WorldNet [99]
Reasoning SPIQA [100], M3Exam [101], Math-Vision [102], MATHCHECK-GEO [103], MathV360K [104], MMSci [105], DUDE [106], CMMMU [107], MathVista [108], ChartingNewTerritories [109], CMMU [110], MindBench [111], MathVerse [112], NPHardEval4V [113], SceMQA [114], CHOPINLLM [115], CharXiv [116], MMWorld [117], MMTab [118], VisualWebBench [119], ChartBench [120], MMC-Benchmark [121], SciGraphQA [122]
II Related Work
II-A MLLM Benchmarks
A variety of benchmark datasets have been proposed to evaluate the capabilities of MLLMs, covering visual perception, language understanding, external knowledge integration, and complex reasoning. To provide a more principled organization, we summarize representative multimodal benchmarks by their primary targeted capability, as shown in Table II. Early datasets such as COCO [123], VQA [124], and Visual7W [125] mainly assess perception and basic question answering abilities. Subsequent benchmarks like GQA [126] and VCR [127] emphasize compositional and commonsense reasoning, while OKVQA [128] and TextVQA [129] introduce challenges involving external knowledge and OCR-based understanding. ScienceQA [130] further targets scientific reasoning in educational contexts. As MLLMs [131, 132, 133] develop, these earlier benchmarks are hard capture their increasingly diverse capabilities. Recent efforts, such as MMBench [38], OmniBench [134], and UniBench [36], aim to provide broader evaluations that include mathematics, medical diagnosis, chart understanding, and instruction following. However, these benchmarks typically rely on heuristic or empirical task categorization, leading to overlapping definitions, inconsistent classification, and limited diagnostic granularity, thus hindering efficiency analysis of model capabilities.
II-B Evaluation for Benchmarks
III Evaluating Benchmark Structures via SEM
III-A Revisiting of Structural Equation Modeling
SEM can be constructed based on causal measurement models [139], organizing multiple observed variables into a coherent structural framework. It consists of two core components: the measurement model [140], which defines the relationships between observed variables and their underlying latent constructs, and the structural model [141], which captures the causal relationships among latent variables.
For evaluating MLLMs capabilities, we focus on the model’s performance across diverse tasks, where task scores collectively determine a latent ability. The specific ability does not preexist to drive task performance, it emerges from the aggregation of task-level outcomes. This paradigm aligns naturally with a formative measurement model, where multiple observed indicators jointly constitute the latent variable. Therefore, we adopt a formative measurement model for benchmark evaluation as shown in Fig. 2. The evaluation framework defines the relationships between observed and latent variables as follows: $X=\Lambda_{x}\xi+\delta$ and $Y=\Lambda_{y}\eta+\epsilon$, where $X$ and $Y$ represent observed exogenous and endogenous variables, $\xi$ and $\eta$ denote latent variables, $\Lambda_{x}$ and $\Lambda_{y}$ are factor loading matrices, and $\delta$ and $\epsilon$ correspond to measurement errors. The structural model captures the causal relationships among latent variables: $\eta=B\eta+\Gamma\xi+\zeta$, where $B$ and $\Gamma$ are path coefficient matrices, and $\zeta$ represents residuals. Together, the measurement and structural models offer a comprehensive framework to assess both direct and indirect effects.
@@ III-A Revisiting of Structural Equation Modeling
III-B Structural Analysis of Existing Benchmarks
@@
Interestingly, we further apply the proposed SEM-based evaluating framework to analyze existing MLLM benchmarks and find that most benchmarks struggle to balance low redundancy with high construct validity. Specifically, we investigate whether the task configurations within current benchmarks provide complementary information while jointly forming meaningful latent constructs.
@@
To illustrate this issue in detail, we conduct a case study on two representative benchmarks: MMBench [38] and BLINK [34]. Leveraging the performance reports of over 190 MLLMs collected via the VLMEvalKit [142] toolkit, we construct formative structural equation models based on the hierarchical task taxonomy specified by each benchmark. In our modeling framework, individual tasks serve as observed indicators, which are grouped into intermediate latent variables representing ability dimensions such as perception, reasoning, and cross-modal understanding. These first-order latent abilities are then aggregated into a second-order latent variable that captures the overall capability of an MLLM. Crucially, to align this latent construct with human-perceived performance, we incorporate the Chatbot Arena [42] score as an external observed indicator of the top-level latent variable. This hierarchical SEM structure enables us to evaluate how well benchmark-defined ability dimensions, as derived from task-level scores, explain human-judged overall performance.
@@
In the case of MMBench, as shown in Fig. 3(a), tasks exhibit high contribution to their corresponding ability dimensions, with outer loadings exceeding 0.85 for most tasks. However, the analysis of Variance Inflation Factor (VIF) values in Fig. 3(b) shows a concerning level of multicollinearity, particularly for tasks like Image Scene (VIF = 13.96) and Attribute Recognition (VIF = 6.12), indicating a high degree of redundancy within the benchmark. This redundancy undermines the overall efficiency of the measurement model, as the tasks are provide insufficiently independent information.
@@
On the other hand, BLINK as shown in Fig. 4 exhibits a different issue. While the VIF values for BLINK are generally low (with a maximum value of 3.09 for Art Style), indicating low task redundancy, the outer loadings of several tasks are lower compared to MMBench. Tasks such as IQ Test (outer loading = 0.354) and Counting (outer loading = 0.686) contribute less to their respective constructs, suggesting that the benchmark lacks sufficiently strong task performance indicators for each cognitive ability dimension.
Despite structural differences, both MMBench and BLINK exhibit limited discriminant validity among their ability dimensions, as indicated by their respective HTMT matrices. Several HTMT values exceed or approach the 0.90 threshold, particularly between conceptually adjacent dimensions, revealing substantial overlap and suggesting that current ability taxonomies in both benchmarks need further conceptual refinement. For instance, in BLINK, the HTMT value between the mid-level and high-level reaches 1.21, highlighting that tasks intended to measure these distinct abilities are highly correlated and may not effectively disentangle the underlying cognitive constructs.
Overall, while MMBench and BLINK differ in their structural characteristics, MMBench showing strong but redundant task contributions, and BLINK presenting low redundancy but weaker indicators. Both suffer from poor discriminant validity in their ability dimensions. These findings expose fundamental limitations in current benchmark design, highlight the need for a more systematic approach to evaluating benchmark structure and the necessity for a principled framework to evaluate the internal structure of MLLM benchmarks and to redefine ability dimensions based on sound cognitive foundations.
III-C SEM-based Evaluation Framework for Benchmark
@@
Dimensional Diversity
it aims to quantify the discriminant validity among latent ability dimensions, thereby characterizing their conceptual independence and potential redundancy. Although existing benchmarks effectively capture multiple facets of MLLMs’ capabilities, overlaps between different dimensions may still exist. To rigorously assess discriminant validity, we adopt the Heterotrait–Monotrait Ratio of correlations (HTMT) [143]. For any two latent dimensions $\xi_{i}$ and $\xi_{j}$, HTMT index is defined as
@@
$D_{\mathrm{div}}=\frac{1}{2*\max_{i\neq j}\mathrm{HTMT}_{ij}},$ (2)
where $\max_{i\neq j}\mathrm{HTMT}{ij}$ represents the maximum HTMT value across all distinct pairs of latent dimensions. The denominator is scaled by 2 to ensure that $D_{\mathrm{div}}\in[0,1]$, given that HTMT values are typically bounded within the empirical range of [0, 1]. A higher $D_{\mathrm{div}}$ indicates stronger discriminant validity and thus greater conceptual distinctiveness among latent constructs.
Task Contribution
it quantifies the relative importance of each task indicator in shaping its corresponding latent construct. In order to capture this notion in a formative measurement context, we adopt the concept of out‑loading [144], which represents the standardized weight of each observed task on the latent dimension. To evaluate the contribution of tasks at the level of the entire benchmark, which consists of multiple ability dimensions $\xi=\{\xi_{1},\dots,\xi_{m}\}$, we define the benchmark-level task contribution score as
$\begin{split}\mathrm{TC}&=\frac{1}{\sum_{i=1}^{n}m_{i}}\sum_{i=1}^{m_{i}}\sum_{j=1}^{n}\left|\lambda_{i,j}\right|,\end{split}$ (3)
the metric $\mathrm{TC}\in[0,1]$ provides a global estimate of task-level informativeness across the entire benchmark, with higher values indicating stronger and more concentrated contributions from individual tasks.
TABLE III: Refined Task Selection Results on MMBench
Capability Representative Task VIF Outer Loading
Attribute Identity Reasoning 1.749 0.893
Reasoning Physical Relation 1.749 0.924
Coarse Image Quality 2.682 0.938
Perception Image Style 2.682 0.954
Cross-instance Action Recognition 1.549 0.818
Perception Spatial Relationship 1.549 0.949
Fine-grained Object Localization 2.113 0.956
Perception OCR 2.113 0.896
Logical Future Prediction 2.793 0.938
Reasoning Structured Image-Text Understanding 2.793 0.959
Relation Physical Property Reasoning 2.133 0.941
Reasoning Social Relation 2.133 0.918
Indicator Validity
it is assessed both substantively and statistically, with the latter focusing on multicollinearity diagnostics among task indicators. For each indicator $x_{i,j}$ associated with latent dimension $\xi_{j}$, we compute the Variance Inflation Factor (VIF) [145] by
@@
where $R^{2}_{x_{i,j}}$ is the coefficient of determination obtained by regressing $x_{i,j}$ on the remaining tasks. A higher VIF indicates stronger collinearity, suggesting that task may not provide unique information beyond what is already captured by others.
To consolidate individual VIF values into a single benchmark-level measure, we define the composite validity score $D_{\mathrm{valid}}$ as the inverse of the geometric mean of all task-level VIFs: